OpenStudy (anonymous):

What is 6^1001 mod 7 @ganeshie8 @Luigi0210 @wio @e.mccormick @shamil98 @dan815 @wolfe8 @nincompoop @Ashleyisakitty @Isaiah.Feynman @kewlgeek555 @tester97 @Mashy @adrynicoleb @sourwing @dpasingh @RaphaelFilgueiras

ganeshie8 (ganeshie8):

\(\large 6^{1001} \equiv (-1)^{1001} \mod 7\)

OpenStudy (tester97):

maybe you can use a ti-84 calculator? anyways im not sure because idk how to solve that

ganeshie8 (ganeshie8):

did u watch amelie movie @tester97

OpenStudy (tester97):

Lol everyone has been asking me that but no its just a screenshot from a clip of the movie.

ganeshie8 (ganeshie8):

its a very good movi... :)

OpenStudy (tester97):

I will check it out ^_^

ganeshie8 (ganeshie8):

il pm u the link on viooz

OpenStudy (tester97):

ok :)

ganeshie8 (ganeshie8):

@TanteAnna it simplifies to -1..

OpenStudy (dan815):

what is this -1^1001 why is it equivalent to that

ganeshie8 (ganeshie8):

\(6 \equiv -1 \mod 7\)

ganeshie8 (ganeshie8):

above expression means : 6 leaves a remainder -1, when divided by 7

OpenStudy (dan815):

oh ok

ganeshie8 (ganeshie8):

another easy way to think it is : (6- (-1)) is divisible by 7

OpenStudy (dan815):

i was thinking 6 mod 7 = 6 36 mod 7 = 1 36*6 mod 7 = 6 36*36 mod 7 = 1 and so on so even exponent is 1 and odd exponent = 6?

ganeshie8 (ganeshie8):

that works !

ganeshie8 (ganeshie8):

leaving a remainder 6 is same as leaving a remainder -1

ganeshie8 (ganeshie8):

remainder = 6 : overflow is 6 remainder = -1 : lacking 1 both are same when talking about remainders

OpenStudy (dan815):

true

ganeshie8 (ganeshie8):

here, -1 makes the life simple as we have a big power

OpenStudy (dan815):

why does it work!

OpenStudy (dan815):

what is 7^100 mod 8?

OpenStudy (dan815):

7?

ganeshie8 (ganeshie8):

it works cuz of below theorem : (can be proven easily) if \(a\) leaves a remainder \(r\) when divided by \(b\), then : \(a^n\) leaves a remainder \(r^n\) when divided by \(b\)

ganeshie8 (ganeshie8):

yup !

OpenStudy (dan815):

wai no even so 1?

ganeshie8 (ganeshie8):

oh yes, (-1)^100 = 1

OpenStudy (dan815):

interesting

ganeshie8 (ganeshie8):

so remainder is 1, not 7

OpenStudy (dan815):

what is 5^100 mod 7 2??

ganeshie8 (ganeshie8):

nope

OpenStudy (dan815):

4?

OpenStudy (dan815):

no no it is 2 WOW O_O how

OpenStudy (dan815):

math is mysterious

ganeshie8 (ganeshie8):

5^100 mod 7 = 2^100 mod 7 = 2^(3*33+1) mod 7 = 2*8^33 mod 7 = 2

OpenStudy (dan815):

56^100 mod 23

ganeshie8 (ganeshie8):

56^100 mod 23 10^100 mod 23

ganeshie8 (ganeshie8):

wat u wana do next uhmm

OpenStudy (dan815):

23-56

ganeshie8 (ganeshie8):

56^100 mod 23 10^100 mod 23 100^50 mod 23 8^50 mod 23

ganeshie8 (ganeshie8):

keep reducing it..

OpenStudy (dan815):

k start from beginning teach me from start

ganeshie8 (ganeshie8):

xD this example ? or congruences ?

OpenStudy (dan815):

congruences

ganeshie8 (ganeshie8):

5 small things we need to cover, to be able to play wid remainders at will

OpenStudy (dan815):